BREAK-EVEN ANALYSIS

The Warren Watch Company sells watches for $27, fixed costs are $175,000, and variable costs are $11 per watch.

  1. What is the firm's gain or loss at sales of 7,000 watches? Enter loss (if any) as negative value. Round your answer to the nearest cent.
    $

    What is the firm's gain or loss at sales of 16,000 watches? Enter loss (if any) as negative value. Round your answer to the nearest cent.
    $
  2. What is the break-even point (unit sales)? Round your answer to the nearest whole number.
    units
  3. What would happen to the break-even point if the selling price was raised to $35?
    -Select-The result is that the break-even point is higher./ The result is that the break-even point remains unchanged./ The result is that the break-even point is lower.
  4. What would happen to the break-even point if the selling price was raised to $35 but variable costs rose to $20 a unit? Round your answer to the nearest whole number.
    -Select-The result is that the break-even point decreases. /The result is that the break-even point remains unchanged./ The result is that the break-even point increases.
Answer #1

Gain = Sales – variable costs – Fixed costs

a.Gain = (27-11)*7,000 – 175,000

= -$63,000

i.e. loss

b.Gain = 16000*16 – 175000

= $81,000

i.e. gain

c.Break even unit sales = Fixed costs/(Selling price per watch – variable cost per watch)

= 175,000/16

= 10,937.5 watches

c. The result is that the break-even point is lower. Since contribution margin will increase

d.The result is that the break even point is higher. since contribution margin will decrease.